OpenStudy (anonymous):

Find the zeros of the equation. 2x^2-x=10+y A. x=2 and x=-5/2 B. x=-2 and x=5/2 C. x=2 and x=-2/5

OpenStudy (anonymous):

Set y equal to zero and use the quadratic formula.

OpenStudy (anonymous):

\[2x^2-x-10=0\] \[ax^2+bx+c\] \[x= \frac{ -b \pm \sqrt(b^2-4ac) }{ 2a }\]

OpenStudy (anonymous):

\[x = \frac{ 1 \pm \sqrt(1^2+4(2)(10)) }{ 2(2) } = -2; \frac{ 5 }{ 2 }\]